OpenStudy (anonymous):

Is 4/7 and 5/8 Equal?

OpenStudy (anonymous):

What have you tried so far?

OpenStudy (anonymous):

Nothing, I for got how to find out if they were equal and i need to know which is bigger.

OpenStudy (anonymous):

\[\frac{4}{7} \stackrel{?}{=}\frac{5}{8}\]\[\implies \frac{4}{7} \cdot 56 \stackrel{?}{=}\frac{5}{8}\cdot 56\]\[\implies 32 \stackrel{?}{=}35\]

OpenStudy (anonymous):

Thanks, but and i remembered the 5th grade way, Butterfly!

OpenStudy (anonymous):

no